OpenStudy (anonymous):

Solve the following system of equations. x + 3y – z = 2 x – 2y + 3z = 7 x + 2y – 5z = –21

OpenStudy (anonymous):

z=16

OpenStudy (anonymous):

for which variable

OpenStudy (anonymous):

y=119

OpenStudy (anonymous):

x=-339

OpenStudy (anonymous):

x+3y-z=2 for y : Move all terms not containing y to the right-hand side of the equation. 3y=-x+z+2 Divide each term in the equation by 3. (3y)/(3)=-(x)/(3)+(z)/(3)+(2)/(3) Simplify the left-hand side of the equation by canceling the common terms. y=-(x)/(3)+(z)/(3)+(2)/(3) Simplify the right-hand side of the equation by simplifying each term. y=(-x+z+2)/(3) for x : x+3y-z=2 Move all terms not containing x to the right-hand side of the equation. x=-3y+z+2 for z: x+3y-z=2 Move all terms not containing z to the right-hand side of the equation. -z=-x-3y+2 Multiply each term in the equation by -1. -z*-1=-x*-1-3y*-1+2*-1 Multiply -z by -1 to get z. z=-x*-1-3y*-1+2*-1 Simplify the right-hand side of the equation by multiplying out all the terms. z=x+3y-2

OpenStudy (anonymous):

x+3y-z=2_x-2y+3z=7_x+2y-5z=-21 Move all terms not containing x to the right-hand side of the equation. x=-3y+z+2_x-2y+3z=7_x+2y-5z=-21 Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is -3y+z+2. x=-3y+z+2_(-3y+z+2)-2y+3z=7_x+2y-5z=-21 Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is -3y+z+2. x=-3y+z+2_(-3y+z+2)-2y+3z=7_(-3y+z+2)+2y-5z=-21 Remove the parentheses around the expression -3y+z+2. x=-3y+z+2_-3y+z+2-2y+3z=7_(-3y+z+2)+2y-5z=-21 Since -3y and -2y are like terms, subtract 2y from -3y to get -5y. x=-3y+z+2_-5y+z+2+3z=7_(-3y+z+2)+2y-5z=-21 Since z and 3z are like terms, add 3z to z to get 4z. x=-3y+z+2_-5y+4z+2=7_(-3y+z+2)+2y-5z=-21 Remove the parentheses around the expression -3y+z+2. x=-3y+z+2_-5y+4z+2=7_-3y+z+2+2y-5z=-21 Since -3y and 2y are like terms, subtract 2y from -3y to get -y. x=-3y+z+2_-5y+4z+2=7_-y+z+2-5z=-21 Since z and -5z are like terms, add -5z to z to get -4z. x=-3y+z+2_-5y+4z+2=7_-y-4z+2=-21 Move all terms not containing y to the right-hand side of the equation. x=-3y+z+2_-5y+4z+2=7_-y=4z-2-21 Subtract 21 from -2 to get -23. x=-3y+z+2_-5y+4z+2=7_-y=4z-23 Multiply each term in the equation by -1. x=-3y+z+2_-5y+4z+2=7_-y*-1=4z*-1-23*-1 Multiply -y by -1 to get y. x=-3y+z+2_-5y+4z+2=7_y=4z*-1-23*-1 Simplify the right-hand side of the equation by multiplying out all the terms. x=-3y+z+2_-5y+4z+2=7_y=-4z+23 Replace all occurrences of y with the solution found by solving the last equation for y. In this case, the value substituted is -4z+23. x=-3y+z+2_-5(-4z+23)+4z+2=7_y=-4z+23 Multiply -5 by each term inside the parentheses. x=-3y+z+2_(20z-115)+4z+2=7_y=-4z+23 Since 20z and 4z are like terms, add 4z to 20z to get 24z. x=-3y+z+2_24z-115+2=7_y=-4z+23 Add 2 to -115 to get -113. x=-3y+z+2_24z-113=7_y=-4z+23 Since -113 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 113 to both sides. x=-3y+z+2_24z=113+7_y=-4z+23 Add 7 to 113 to get 120. x=-3y+z+2_24z=120_y=-4z+23 Divide each term in the equation by 24. x=-3y+z+2_(24z)/(24)=(120)/(24)_y=-4z+23 Simplify the left-hand side of the equation by canceling the common terms. x=-3y+z+2_z=(120)/(24)_y=-4z+23 Simplify the right-hand side of the equation by simplifying each term. x=-3y+z+2_z=5_y=-4z+23 Replace all occurrences of z with the solution found by solving the last equation for z. In this case, the value substituted is 5. x=-3y+(5)+2_z=5_y=-4z+23 Replace all occurrences of z with the solution found by solving the last equation for z. In this case, the value substituted is 5. x=-3y+(5)+2_z=5_y=-4(5)+23 Remove the parentheses around the expression 5. x=-3y+5+2_z=5_y=-4(5)+23 Add 2 to 5 to get 7. x=-3y+7_z=5_y=-4(5)+23 Multiply -4 by each term inside the parentheses. x=-3y+7_z=5_y=-20+23 Add 23 to -20 to get 3. x=-3y+7_z=5_y=3 Replace all occurrences of y with the solution found by solving the last equation for y. In this case, the value substituted is 3. x=-3(3)+7_z=5_y=3 Multiply -3 by each term inside the parentheses. x=-9+7_z=5_y=3 Add 7 to -9 to get -2. x=-2_z=5_y=3 This is the solution to the system of equations. x=-2_z=5_y=3